OpenStudy (anonymous):

help pls calculate the amount of acetic acid and acetate needed to prepare the acetate buffer system, two liters of 5 mM buffer solution with pH 5.2. pKa of acetate buffer is 4.76.

OpenStudy (aaronq):

so start with the Henderson Hasselbalch equation \(pH=pKa+ log\dfrac{[AcCOO^-]}{[AcCOOH]}\) plug your values in and find the \(\dfrac{[AcCOO^-]}{[AcCOOH]}\) ratio. From there you can find the necessary amounts (moles) in your 2 L system.

OpenStudy (anonymous):

@aaronq does the ratio only apply for 1 L?

OpenStudy (aaronq):

well the ratio is in terms of concentration, so you wanna find the total moles first, then find how much there is of each with the ratio.

OpenStudy (anonymous):

im confused

OpenStudy (aaronq):

whats the total amount of moles of the buffer?

OpenStudy (anonymous):

0.005 M?

OpenStudy (aaronq):

that's the molarity, what about moles?

OpenStudy (anonymous):

@aaronq how to get the moles out from that?

OpenStudy (aaronq):

use the molarity formula and the volume of the solution